OpenStudy (anonymous):

Triangles ABC and DEF. Coordinates for A are 0, 2; for B 2, 4; and for C 1, negative 2. Given: ΔABC ≅ ΔEFD What is the length of Segment line FE rounded to the nearest tenth?

OpenStudy (anonymous):

2.8 3.4 4.1 I think its this one but im not sure ?? 6.1

OpenStudy (anonymous):

Find the distance between point A and B

OpenStudy (anonymous):

answer is 2.8

OpenStudy (paxpolaris):

\[FE = AB = \sqrt{(4-2)^2+(2-0)^2}= \sqrt{2)^2+2^2}=\sqrt8\]